### v1

@@ 5.3 Effectiveness of two-stage training strategy
Fig. 11.: Average Recall@1 results of LifelongPR with different training strategies.
5.4 Parameter analysis
Table 6 presents the CL results of LifelongPR with different temperature parameters. The temperature parameter $\tau$ serves as a key hyperparameter in replay sample selection, where larger values lead to more uniform allocation of replay samples. Experimental results demonstrate that our method achieves optimal performance when $\tau$ is set to 4.0, which is therefore recommended as the default parameter setting.
Table 6: Continual learning results of LifelongPR with different temperature thresholds $\tau$. Note: The prompt module is not employed.
Temperature ($\tau$) mIR@1(%$\uparrow$ mR@1(%$\uparrow$ F(%$\downarrow$
@@ 5.4 Parameter analysis
5.5 Deficiencies and future work
While our method demonstrates promising performance, several limitations remain. Fig. 5.5 presents three representative failure cases. Case 1 involves a T-junction, while Case 3 exhibits insufficient scene characteristics in the query submap, where both the retrieved top 1 submaps and the query submap show high similarity. Case 2 represents a typical street scene with trees lining the roadsides, lacking distinctive geographical features. Among these cases, Cases 1 and 2 could potentially be addressed by enhancing network generalization and anti-forgetting capabilities. However, Case 3 involves repetitive scenes that cannot be resolved through single-shot retrieval. Furthermore, the proposed prompt module also exhibits catastrophic forgetting, and the CL approach built upon it necessitates that the backbone network must be initially trained to ensure robust feature extraction capability and strong generalizability. In the future, we plan to incorporate theories and approaches of domain generalization (Zhou et al., 2022) and contrastive learning (Liu et al., 2021) to enhance the backbone network’s feature extraction capability and generalizability, thereby fundamentally improving the anti-forgetting capability of PCPR models. Furthermore, we will research on online CL for PCPR, enabling the model to evolve incrementally during runtime operation (Aljundi et al., 2019).
Fig. 12.: Three bad cases of LifelongPR trained on Seq1 and Seq2. Column 1: query; Column 2: top 1 submaps retrieved by the model trained on Oxford; Column 3: top 1 submaps retrieved by the model trained on Seq1; Column 4: top 1 submaps retrieved by the model trained on Seq2.
6 Conclusion
@@
[…]
7 Declaration of interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ V-C Effectiveness of two-stage training strategy
Figure 12: Average Recall@1 results of LifelongPR with different training strategies. (a) no prompt, (b) one-stage, (c) prompt only, (d) two-stage (ours).
V-D Parameter analysis
Table VI presents the CL results of LifelongPR with different temperature parameters. The temperature parameter $\tau$ serves as a key hyperparameter in replay sample selection, where larger values lead to more uniform allocation of replay samples. Experimental results demonstrate that our method achieves optimal performance when $\tau$ is set to 4.0, which is therefore recommended as the default parameter setting.
TABLE VI: Continual learning results of LifelongPR with different temperature thresholds $\tau$. Note: The prompt module is not employed.
[…]
@@ V-D Parameter analysis
V-E Deficiencies and future work
While our method demonstrates promising performance, several limitations remain. Fig. 13 presents three representative failure cases. Case 1 involves a T-junction, while Case 3 exhibits insufficient scene characteristics in the query submap, where both the retrieved top 1 submaps and the query submap show high similarity. Case 2 represents a typical street scene with trees lining the roadsides, lacking distinctive geographical features. Among these cases, Cases 1 and 2 could potentially be addressed by enhancing network generalization and anti-forgetting capabilities. However, Case 3 involves repetitive scenes that cannot be resolved through single-shot retrieval. Furthermore, the proposed prompt module also exhibits catastrophic forgetting, and the CL approach built upon it necessitates that the backbone network must be initially trained to ensure robust feature extraction capability and strong generalizability. In the future, we plan to incorporate theories and approaches of domain generalization [59] and contrastive learning [60] to enhance the backbone network’s feature extraction capability and generalizability, thereby fundamentally improving the anti-forgetting capability of PCPR models. Furthermore, we will research on online CL for PCPR, enabling the model to evolve incrementally during runtime operation [61].
Figure 13: Three bad cases of LifelongPR trained on Seq1 and Seq2. Column 1: query; Column 2: top 1 submaps retrieved by the model trained on Oxford; Column 3: top 1 submaps retrieved by the model trained on Seq1; Column 4: top 1 submaps retrieved by the model trained on Seq2.
VI Conclusion
@@
Biography Section
[Uncaptioned image] Xianghong Zou received the B.S. degree in geomatics engineering, the M.S. degree, and the Ph.D. degree in photogrammetry and remote sensing from Wuhan University, Wuhan, China, in 2016, 2019, and 2024, respectively. He is currently a post-doctoral researcher with the School of Advanced Manufacturing, Nanchang University, Nanchang, China. His research interests include point cloud data processing, global localization, and 3D change detection.
[Uncaptioned image] Jianping Li received the B.S. degree in GIS and the Ph.D. degree in photogrammetry and remote sensing from Wuhan University, Wuhan, China, in 2015 and 2021, respectively. He is currently a Research Fellow with the School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore. His research interests include 3D sensing system integration, UAV/UGV mapping, robot perception, and point cloud data processing.
[Uncaptioned image] Zhe Chen received the B.S. degree from the School of Earth Sciences and Engineering, Hohai University, Nanjing, China, in 2020. He is currently pursuing the Ph.D. degree with the State Key Laboratory of Information Engineering in Surveying, Mapping, and Remote Sensing, Wuhan University, Wuhan, China. His research interests include image/point cloud processing and their applications in urban morphology.
[Uncaptioned image] Zhen Cao is a first-year Ph.D. student at the State Key Laboratory of Information Engineering in Surveying, Mapping and Remote Sensing (LIESMARS), Wuhan University, under the supervision of Prof. Zhen Dong and Prof. Bisheng Yang. He obtained his B.Eng. degree in Photogrammetry and Remote Sensing from the School of Geodesy and Geomatics (SGG), Wuhan University. His research interest lies in the field of 3D Computer Vision, particularly including point cloud completion, scene understanding, and intelligent transportation systems (ITS).
[Uncaptioned image] Zhen Dong received the B.E. and Ph.D. degrees in remote sensing and photogrammetry from Wuhan University, in 2011 and 2018, respectively. He is currently a Professor of the State Key Laboratory of Information Engineering in Surveying, Mapping and Remote Sensing (LIESMARS), Wuhan University. His research interest lies in the field of 3D Computer Vision, particularly including 3D reconstruction, scene understanding, point cloud processing as well as their applications in the intelligent transportation systems, digital twin cities, urban sustainable development and robotics.
[Uncaptioned image] Qiegen Liu received the Ph.D. degree in biomedical engineering from Shanghai Jiaotong University, Shanghai, China, in 2012. During his time, he has been a Guest Speaker with the Institute of Computational Biology, Chinese Academy of Sciences (CAS), Beijing, China, and the Laubert Center for Medical Imaging, Shenzhen Institute of Advanced Technology (SIAT), Shenzhen, China. He did his Postdoctoral Work with UIUC, Champaign, IL, USA, in 2015, and the University of Calgary, Calgary, AB, Canada, in 2016. He is mainly engaged in sparse and deep learning representations and their applications in medical imaging and image processing.
[Uncaptioned image] Bisheng Yang received the B.S. degree in engineering survey, the M.S. degree, and the Ph.D. degree in photogrammetry and remote sensing from Wuhan University, China, in 1996, 1999, and 2002, respectively. From 2002 to 2006, he held a post-doctoral position at the University of Zurich, Switzerland. Since 2007, he has been a Professor with the State Key Laboratory of Information Engineering in Surveying, Mapping and Remote Sensing (LIESMARS), Wuhan University, where he is currently the Director of LIESMARS. His main research interests comprise 3-D geographic information systems, urban modeling, and digital city. He was a Guest Editor of the ISPRS Journal of Photogrammetry and Remote Sensing, and Computers & Geosciences.